OpenStudy (anonymous):

Integrate sqrt(1+sqrt(x))/x

OpenStudy (lgbasallote):

\[\huge \int \frac{\sqrt{1 + \sqrt x}}x\] right?

OpenStudy (anonymous):

yep

OpenStudy (lgbasallote):

have you tried substitution?

OpenStudy (anonymous):

I'm not sure what the best sub. option would be. I believe I've faced a similar problem, I just can't recall.....

OpenStudy (lgbasallote):

let a = \(\sqrt x\) square both sides\[a^2 = x\] derive both sides \[2ada = dx\] so your integral will become \[\huge \int \frac{\sqrt{1 - a}}{a^2} (2ada)\]

OpenStudy (experimentx):

try sqrt(1 + sqrt(x)) = y ... better.